OpenStudy (anonymous):

If f'(x) =(2-x)^2 and the y-intercept of f(x) is 4/3, find the rule for f(x)

OpenStudy (anonymous):

Do we have to integrate f'(x)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So from the outside it will be (....)3/3

OpenStudy (anonymous):

\[f \left( x \right)=\int\limits \left( 2-x \right)^2 ~dx=\frac{ \left( 2-x \right)^3 }{ 3 }(-1)+c\] when x=0,f(x)=4/3 find c